Metal oxide have 72.4% and70% metal second oxide of metal is M2O3 find first

Grade:11

1 Answers

Ishika dammani
6 years ago
Ratio of metal and oxygen in second oxide, M2O3 = 70:30Let molecular mass of metal = MTherefore, the percentage by weight of the metal in the oxide(2×M×100)/(2×m+3×16)=70200M=140M+336060M=3360M=56kg Now,Moles of metal in first oxide = 72.4 / 56 = 1.29Moles of oxygen in second oxide = 27.6 / 16 = 1.72Ration of moles of metal and oxygen = 1.29 : 1.72 = 1 : 1.33 = 3 : 4Hence, formula of the first oxide = M3O4